OpenStudy (anonymous):

A block was unable to slide down a rough , heavy wedge of angle 30 with the horizontal the coefficient of friction is sqrt3/2 the min horizontal accl to be given to the wedge so the block moves

OpenStudy (anonymous):

*I know nothing of how to use friction, but bookmark anyway.

OpenStudy (anonymous):

i think Pseudo Force has something to play here

OpenStudy (experimentx):

try this 

OpenStudy (experimentx):

hold on!!

OpenStudy (anonymous):

Mg sin30 - Umg cos 30 = ma **** we have to find the min acc of wedge so that the mass moves

OpenStudy (experimentx):

what's the value of this? Mg sin30 - Umg cos 30 =

OpenStudy (experimentx):

sin(30) = sqrt(3/2) cos(30) = ?

OpenStudy (experimentx):

sin(30) - sqrt(3/2) cos(30) = ?

OpenStudy (anonymous):

thats hw i approched...this...but nw i came to knw that....we have to find the accl of wedge

OpenStudy (ujjwal):